OpenStudy (anonymous):

Perform the operation and write the result in standard form. (10 + 9i) - (-5 - i)

OpenStudy (anonymous):

somebody please help , thanks =)

OpenStudy (anonymous):

everybody feel free to help the more the marrior lol =)

OpenStudy (anonymous):

it's like combining like terms.. combine the real parts and combine the imaginary parts... real part: 10 - (-5) = 15 imaginary part: 9i - (-i) = 10i so you have: 15 + 10i

OpenStudy (anonymous):

i got ...... 15+8i ???

OpenStudy (kainui):

remember the negative on the second term turns -(-5-i) into (5+i) not (5-i)! Almost right leah!

OpenStudy (anonymous):

o, 5+8i ????

OpenStudy (kainui):

Oh nope, backwards. It changes into (10+9i)+(5+i) and when you drop the parenthesis you get 10+9i+5+i

OpenStudy (asnaseer):

try doing this one step at a time. so starting with this: (10 + 9i) - (-5 - i) we first remove the braces around (-5 - i) to get: (10 + 9i) + 5 + i [because - & - = +]

OpenStudy (asnaseer):

understand so far?

OpenStudy (asnaseer):

have you heard of the rule that two minuses make a plus?

OpenStudy (anonymous):

15+10i ???

OpenStudy (asnaseer):

yes - that is correct